OpenStudy (anonymous):

What are the 4 steps for graphing the quadratic function given. (Use ^ on the keyboard to indicate an exponent.) y = -x^2 - 4x - 3

OpenStudy (anonymous):

I solved part of this, I know the line of symmetry is x=-2, and I think the vertex is 1,

OpenStudy (anonymous):

I need help making the table chart. I can graph this I just need help solving to get the equations for the graph.

OpenStudy (skullpatrol):

Do you know how to set up a chart?

OpenStudy (anonymous):

Yeah I just have no idea where to start!

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

Start at the vertex :-)

OpenStudy (anonymous):

I have no idea what's after that, so am I suppose to do -3 next?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

so it would be ? \[-(-3)^2 - 4(-3) -3 = -6 + 12 - 3 =1 ?\]

OpenStudy (skullpatrol):

-(-3)^2 = -9 not -6

OpenStudy (anonymous):

Ahhh I knew it! ugh I deleted it!

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

@skullpatrol so then under the "X" part of the chart do i also put -4, and then -1 below -2, and then 0?

OpenStudy (skullpatrol):